# On computational complexity and average-case hardness of shallow-depth boson sampling

### v1

@@ VIII Concluding remarks
1. Our result demonstrates the average-case hardness for additive imprecision $2^{-O(N^{\gamma+1}(\log N)^{2})}$. Indeed, there still remains a gap to the desired additive imprecision for the simulation hardness $2^{-(\gamma-1)N\log N-O(N)}$ in Theorem 5. Hence, closing this gap would be an ultimate challenge to the full achievement of classical intractability; more advanced proof techniques are required to reduce this gap. Here, one can take the following approach: finite-size numerical experiments suggest that the output distributions of local random circuits in the butterfly circuit architecture (Definition 2) are close enough to those of global Haar random circuits [37]. Accordingly, if one can analytically prove that the distance between those output distributions is close enough, we can directly obtain a better imprecision level $2^{-O(N\log N)}$ by results in [21, 3, 22], which employed the global Haar random circuits.
@@
Another possible approach for reducing the imprecision gap is to perturb a random circuit matrix in a different way from the Cayley transform (Definition 6), i.e., as depicted in [22]. Specifically, instead of perturbing each random gate, one can perturb a submatrix $X$ of our random circuit matrix $U$ with a worst-case matrix $A$ as $X(\theta)=(1-\theta)X+\theta A$ for $\theta\in[0,1]$. Here, a degree of polynomial $|\text{Per}(X(\theta))|^{2}$ is $2N$, which is lower than ours derived by the Cayley transform. Therefore, if one can prove that $X(\theta)$ is distributed similarly to $X$ for small $\theta$, we expect that we can also obtain a better imprecision level by using the same interpolation method. However, the above approach requires one to figure out a global circuit distribution generated by the convolution of local circuit distributions. Although we believe that this problem can be resolved using techniques from random matrix theories, we have not yet developed a complete analysis. Hence, we leave it as an open question.
@@
2. Another important challenge that should be addressed is to find the classical simulation hardness of noisy boson sampling, for general types of physical noise beyond the photon loss noise model we have dealt with so far. To do so, as described in [41, 21, 3], employing the threshold theorem would be a viable choice for this goal. Specifically, the threshold theorem for general types of noise in boson sampling setups has to be developed. This requires an efficient error detection code for general types of error using linear optical elements, for any multi-mode Fock state or Gaussian state input. However, to the best of our knowledge, such an error detection code does not exist. Hence, constructing this error detection code would be a crucial step toward the hardness of noisy boson sampling, which will contribute to a more noise-tolerant demonstration of quantum advantage with boson sampling. We leave this problem as another open question.
Acknowledgement
@@
We thank Bill Fefferman for insightful discussions. The authors acknowledge support from the National Research Foundation of Korea (NRF) grants funded by the Korean government (Grant Nos. NRF-2023R1A2C1006115, NRF-2022M3K4A1097117, and NRF-2022M3E4A1076099) via the Institute of Applied Physics at Seoul National University, the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-2021-0-01059 and IITP-20232020-0-01606). B.G. was also supported by the education and training program of the Quantum Information Research Support Center, funded through the National Research Foundation of Korea (NRF) by the Ministry of Science and ICT (MSIT) of the Korean government (No.2021M3H3A1036573).
@@
Appendix A Previous foundations: Average-case hardness of boson sampling
@@
In this appendix, we argue the existing proof technique employed for the simulation hardness of boson sampling, specifically in the context of the approximate simulation within total variation distance error [1]. The current state-of-the-art proof technique for the hardness of sampling problems like boson sampling essentially builds upon Stockmeyer’s algorithm about approximate counting [38]. Specifically, given a classical sampler that outputs a sample from a given output distribution, Stockmeyer’s algorithm enables one to multiplicatively estimate a fixed output probability of the sampler, within complexity class $\rm{BPP}^{\rm{NP}}$.
@@
Moreover, by choosing random circuit instances that have symmetry over the outcomes, one can reduce the average-case instances for the hardness from outcome instances to circuit instances, which is called the hiding property. For the boson sampling case, global Haar random unitary (i.e., unitary matrix drawn from Haar measure on U(M), for mode number M) satisfies this condition. In detail, instead of randomly choosing the outcome, we can fix the outcome by applying a random permutation to the global Haar random unitary distribution, which is still Haar distributed from its symmetric property. This hiding property plays an important role in the current proofs of the average-case hardness, as it enables one to establish worst-to-average-case reduction. Specifically, as the output probability of boson sampling can be written as a low-degree polynomial of input circuit (matrix) values, it allows one to infer the value of a worst-case instance from the output probability of many average-case circuit instances. Hence, the average-case hardness argument for boson sampling is typically used in this context, i.e., average-case hardness over random circuit instances, for a fixed outcome [1, 3, 21].
@@
Accordingly, the crucial step for the classical simulation hardness of approximate sampling is to prove the average-case hardness for the desired imprecision level. While there have been many impressive results about the average-case hardness of boson sampling [1, 3, 21, 22], the average-case hardness for the desired imprecision level is not yet fully demonstrated. Still, there exists a gap between the imprecision level of average-case hardness in the strongest existing results and the imprecision level of average-case approximation problem. Hence, closing this imprecision gap remains the ultimate challenge for the fully theoretically guaranteed computational advantage of approximate boson sampling.
Appendix B Proof of Theorem 3
The proof of Theorem 3 can be greatly simplified by introducing the following two Lemmas.
@@ Appendix B Proof of Theorem 3
For an arbitrary given poly-sized $n$-qubit quantum circuit $Q$, there exists a constant depth linear optical circuit $C$ such that for $M\geq 2N$ and $N=\text{poly}(n)$,
[…]
where $c_{Q}$ is a $Q$ dependent constant which can be efficiently computed, $\hat{\mathcal{U}}(C)$ is a unitary operator corresponding to the circuit $C$, and |$\ket{I}$ is an $M$-mode Fock state composed of $N$ single photon states and vacuum states for the rest modes.
@@
We revise the results by [25], for a more rigorous analysis of the allowed additive imprecision level for the worst-case hardness of shallow-depth boson sampling. Ref. [25] proposed that certain 4-depth linear-optical circuits with post-selection can simulate universal quantum computing. Specifically, to simulate any poly-sized quantum circuit $Q$ on $n$ qubits, there exists a measurement-based quantum computation (MBQC) scheme using constant depth brickwork graph state of maximally poly($n$) qubits [42, 43, 44]. The corresponding scheme can also be implemented in a linear optical system via KLM scheme [45] with post-selection, using $N=\text{poly}(n)$ number of single photon states over $M\geq 2N$ modes (requirements for the dual-rail encoding), which can simulate the quantum circuit composed of $O(N)$ number of gates.
Therefore, given the circuit in [25] and an appropriate dual-rail encoded state |$\ket{I}$, the output probability of any quantum circuit $Q$ can be represented as Eq. (21), where $c_{Q}$ denotes the product of post-selection probabilities for gate implementations. To compute $c_{Q}$, we need to figure out the required number of gates to implement the circuit $Q$ and their probabilities to be post-selected. More precisely, from [25], post-selection occurs for two cases: $(i)$ for the CZ gate to implement the brickwork graph state, and $(ii)$ for the gate set {CX, T, H} (i.e., universal set of gates) which can be implemented by measurement of the graph state.
Hence, $c_{Q}$ can be expressed as
@@
where $p_{k}$ denotes post-selection probability of $k$ gate (e.g., $p_{\text{CZ}}$ is $2/27$ in [46]), and $\Gamma_{k}$ denotes the number of $k$ gate to implement the circuit $Q$. By counting the number of each gate to implement the circuit $Q$, $c_{Q}$ can be computed efficiently. ∎

### v2

@@ VIII Concluding remarks
1. Our result demonstrates the average-case hardness for additive imprecision $2^{-O(N^{\gamma+1}(\log N)^{2})}$. Indeed, there still remains a gap to the desired additive imprecision for the simulation hardness $2^{-(\gamma-1)N\log N-O(N)}$ in Theorem 5. Hence, closing this gap would be an ultimate challenge to the full achievement of classical intractability; more advanced proof techniques are required to reduce this gap. Here, one can take the following approach: finite-size numerical experiments suggest that the output distributions of local random circuits in our circuit architecture (i.e., Definition 2) are close enough to those of global Haar random circuits (see, e.g., Ref. [46] and Appendix H). Accordingly, if one can analytically prove that the distance between those output distributions is close enough, we can directly obtain a better imprecision level $2^{-O(N\log N)}$ by results in [28, 3, 29, 30], which employed the global Haar random circuits.
@@
Another possible approach for reducing the imprecision gap is to perturb a random circuit matrix in a different way from the Cayley transform (i.e., Definition 8), i.e., as depicted in [29]. Specifically, instead of perturbing each random gate, one can perturb a submatrix $X$ of our random circuit matrix $U$ with a worst-case matrix $A$ as $X(\theta)=(1-\theta)X+\theta A$ for $\theta\in[0,1]$. Here, a degree of polynomial $|\text{Per}(X(\theta))|^{2}$ is $2N$, which is lower than ours derived by the Cayley transform. Therefore, if one can prove that $X(\theta)$ is distributed similarly to $X$ for small $\theta$, we expect that we can also obtain a better imprecision level by using the same interpolation method. However, the above approach requires one to figure out a global circuit distribution generated by the convolution of local circuit distributions. Although we believe that this problem can be resolved using techniques from random matrix theories, we have not yet developed a complete analysis. Hence, we leave it as an open question.
@@
2. Another important challenge that should be addressed is the extension of the result in Sec. VII, i.e., finding the classical simulation hardness of noisy Gaussian Boson Sampling with photon loss, or Boson Sampling for more general types of noise. As described in [50, 28, 3], employing the threshold theorem would be a viable choice for this goal, which requires efficient error detection codes using only linear optical elements for these setups. However, to the best of our knowledge, such an error detection code does not yet exist. That is, in contrast to the Boson Sampling case, we cannot post-select the so-called ‘no loss’ event in the Gaussian Boson Sampling case because the input photon number is not fixed in this case. This is also for the Boson Sampling with general types of noise because we cannot post-select no error event by using only photon number measurements. Hence, constructing any of these error detection codes would be a crucial step toward the classical hardness of noisy (Gaussian) Boson Sampling, which will contribute to a more noise-tolerant demonstration of quantum advantage with (Gaussian) Boson Sampling. We leave this problem as another open question.
3. One of the existing gaps between theoretical Boson Sampling hardness arguments and Boson Sampling experiments to date is the ratio between mode number $M$ and photon number $N$. Generally, most of the current theoretical Boson Sampling hardness arguments (including our current work) consider only collision-free outcomes and rule out collision outcomes due to the convenience of mathematical analysis [1, 3, 27, 28, 30]. This calls for the “dilute” regime $M=\Omega(N^{2})$ to ensure the dominance of collision-free outcomes over the probability weight. In contrast, all the experiments to date [4, 5, 6, 7] operate in the “saturated regime” where the number of modes is comparable to the number of photons. Therefore, to reduce this gap, extending our hardness result of shallow-depth Boson Sampling to the saturated regime $M=O(N^{\gamma})$ for $21\leq\gamma<$ 2 is crucial, similarly to [29] for the general Boson Sampling case. To do so, one also needs to consider collision outcomes in the analysis because collision outcomes now dominate the probability weight in this regime. Here, given that our worst-case hardness result in Theorem 3 holds for a linear number of modes $M=\Omega(N)$, this opens the possibility of establishing average-case hardness in the saturated regime by extending the worst-to-average-case reduction from Theorem 3 to include collision outcomes. However, since the collision outcomes do not interchange with each other generally by permutations (in contrast to collision-free outcomes), our proof technique in Sec. IV using random permutation does not work for general collision outcomes. Therefore, more advanced proof techniques are required to address this challenge. In Appendix G, we present an approach that offers useful insight toward extending our average-case hardness result to include collision outcomes. Also, another possible way is to analytically prove that the local random circuit for our circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ is close enough to those of global Haar random circuits (numerically shown in [46] and Appendix H). If this can be done, we can employ the previous argument [29] that used the global Haar random circuits, for our circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ in the saturated regime.
4. We finally remark that if we only use the local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ for our shallow-depth circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ as defined in Definition 4, the bosonic birthday paradox is not guaranteed because of the absence of translation invariance property (in contrast to the global Haar random unitary circuit). Hence, for the sake of analytical completeness, we introduced the additional random permutation circuit at the input to guarantee the bosonic birthday paradox and suppress the occurrence of collision outcomes. In fact, from an experimental perspective, this setup can be realized by constructing the local random circuit $\mathcal{H}_{\mathcal{A}}$ and preparing a random input configuration for each experiment. Furthermore, if one can analytically verify that the local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ has a permutation invariance property for $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$, then we can safely leave out this random permutation circuit part, thereby simplifying the average-case circuits and reducing the required circuit depth for the average-case hardness. We leave in Appendix H numerical analysis for the bosonic birthday paradox for our local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$, which gives evidence that only the local random circuit ensemble might be sufficient for the average-case hardness result.
Acknowledgements
@@
We thank Bill Fefferman for insightful discussions. This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (Nos. RS-2024-00413957, RS-2024-00438415, NRF-2023R1A2C1006115, and RS-2024-00437191) and by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (IITP-2025-RS-2020-II201606). B.G. was supported by the education and training program of the Quantum Information Research Support Center, funded through the National Research Foundation of Korea (NRF) by the Ministry of Science and ICT (MSIT) of the Korean government (No.2021M3H3A1036573). C.O. was supported by the National Research Foundation of Korea (NRF) Grants (No. RS-2024-00431768 and No. RS-2025-00515456) funded by the Korean government (Ministry of Science and ICT (MSIT)).
@@
Appendix A Previous foundations on average-case hardness of Boson Sampling
@@
In this appendix, we introduce the existing proof technique used to establish the classical hardness of Boson Sampling, specifically in the context of approximate simulation within a bounded total variation distance error. [1]. The current state-of-the-art proof technique for the hardness of sampling problems like Boson Sampling essentially builds upon Stockmeyer’s algorithm about approximate counting [47]. Specifically, given a classical sampler that outputs a sample from a given output distribution, Stockmeyer’s algorithm enables one to multiplicatively estimate a fixed output probability of the sampler, within complexity class $\rm{BPP}^{\rm{NP}}$.
@@
Moreover, by choosing random circuit instances that have symmetry over the outcomes, one can reduce the average-case instances for the hardness from outcome instances to circuit instances, which is called the hiding property. For the Boson Sampling case, global Haar random unitary (i.e., unitary matrix drawn from Haar measure on U(M), for mode number M) satisfies this condition. In detail, instead of randomly choosing the outcome, we can fix the outcome by applying a random permutation to the global Haar random unitary distribution, which is still Haar distributed due to its translation invariance property. This hiding property plays an important role in the current proofs of the average-case hardness, as it enables one to establish worst-to-average-case reduction in terms of circuit instances. Specifically, as the output probability of Boson Sampling can be written as a low-degree polynomial of input circuit (matrix) values, it allows one to infer the value of a worst-case instance from the output probability of many average-case circuit instances. Hence, the average-case hardness argument for Boson Sampling is typically used in this context, i.e., average-case hardness over random circuit instances, for a fixed outcome [1, 3, 28].
@@
Accordingly, the crucial step for the classical simulation hardness of approximate sampling is to prove the average-case hardness for the desired imprecision level. While there have been many impressive results and technical developments about the average-case hardness of Boson Sampling [1, 3, 27, 28, 29, 30], the average-case hardness for the desired imprecision level is not yet fully demonstrated. Still, there exists a gap between the imprecision level of average-case hardness in the strongest existing results and the imprecision level of average-case approximation problem. Hence, closing this imprecision gap remains the ultimate challenge for the complete theoretical demonstration of the classical hardness of approximate Boson Sampling.
Appendix B The bosonic birthday paradox for the shallow average-case circuits
Throughout our work, we consider only collision-free outcomes in our average-case hardness arguments instead of all possible outcomes. In this appendix, we examine whether the bosonic birthday paradox [1] holds for our average-case circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$. More specifically, we investigate whether collision-free outcomes are observed with high probability in the so-called “dilute” regime, where the number of modes satisfies $M=\Omega(N^{2})$.
To claim the classical hardness and quantum computational advantage of Boson Sampling based solely on collision-free outcomes, it must be ensured that these collision-free outcomes occupy at least a not-a-very-small portion of the probability weight over all possible outcomes. To understand why, let us denote the probability weight corresponding to the collision-free outcomes by $\varepsilon$. For the probability weight of collision-free outcomes $\varepsilon$, the “adversarial” classical sampler that outputs only collision outcomes is also a good sampler, if the allowed total variation distance error for the classical sampler is larger than $\varepsilon$. Using this adversarial sampler combined with Stockmeyer’s algorithm that multiplicative approximates the output probability of the sampler, the obtained estimates of output probabilities of collision-free outcomes are always “0”, since this sampler does not output collision-free outcomes. Therefore, even if we derive the average-case #P-hardness of estimating output probabilities for collision-free outcomes, we cannot solve this #P-hard problem by Stockmeyer’s reduction if the given sampler is adversarial; thereby, we cannot derive the classical hardness result of sampling (which allows total variation distance error) from the average-case #P-hardness.
Accordingly, to avoid this issue and claim the classical hardness result using only collision-free outcomes, the bosonic birthday paradox is crucial, which suppresses the occurrence of collision outcomes over the experiments. This property holds for global Haar random circuits in the dilute regime $M=\Omega(N^{2})$ [1], but not guaranteed for any other random circuit ensemble, including our circuit ensemble. In fact, collision-free outcomes dominate the outcome space in the regime $M=\Omega(N^{2})$ since the portion of collision-free outcomes over all possible outcomes can be bounded as $\binom{M}{N}/\binom{M+N-1}{N}>1-\frac{N^{2}}{M}$. Accordingly, one might expect that the bosonic birthday paradox automatically holds for any circuit ensemble. However, even in this regime, the probability weight might be highly concentrated on collision outcomes for some circuit ensembles (i.e., anticoncentration is not guaranteed for arbitrary circuit ensembles), such that the probability weight of the collision-free outcomes is still too small. Hence, it is crucial to show that the bosonic birthday paradox for our circuit ensemble, such that our circuit ensemble outputs collision-free outcomes with high probability.
Based on this understanding, we now argue the bosonic birthday paradox for the circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ in the dilute regime $M=\Omega(N^{2})$. We first review a lemma given in Ref. [1, Lemma C.1], which is indeed a crucial ingredient for the bosonic birthday paradox.
Lemma 3 (Unitary Pigeonhole Principle [1]).
Partition a finite set $[K]$ into a “good part” $G$ and “bad part” $B=[K]\setminus G$. Also, let $\hat{U}$ be an arbitrary $K$ by $K$ unitary matrix. Suppose we choose an element $x\in G$ uniformly at random, apply $\hat{U}$ to |$\ket{x}$, and then measure $\hat{U}\ket{x}$ in the standard basis. Then, for measurement outcome $y$, we have $\Pr[y\in B]\leq|B|/|G|$.
Here, let us consider the finite set $[K]$ in the lemma as a set of all possible outcomes of Boson Sampling for $N$ photons and $M$ modes, i.e., the set of possible $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in[N]$ with the constraint $\sum_{i=1}^{M}s_{i}=N$ (such that, $|K|=\binom{M+N-1}{N}$). Also, we consider $\hat{U}$ in the lemma as a unitary evolution over the Hilbert space of $N$ photons and $M$ modes, characterized by a given unitary circuit of Boson Sampling.
We now consider the good part $G$ in the lemma as a set of collision-free $\bm{s}=(s_{1},\dots,s_{M})$, such that each $s_{i}\in\{0,1\}$ instead of $s_{i}\in[N]$, which we will denote as $G_{M,N}$. Likewise, let $B_{M,N}$ be the bad part in the lemma, i.e., a set of collision $\bm{s}$. Then, observe that $|G_{M,N}|=\binom{M}{N}$, $|B_{M,N}|=\binom{M+N-1}{N}-\binom{M}{N}$, and thus we have
$\frac{|B_{M,N}|}{|G_{M,N}|}=\frac{\binom{M+N-1}{N}-\binom{M}{N}}{\binom{M}{N}}<\frac{\frac{N^{2}}{M}}{1-\frac{N^{2}}{M}}\leq\frac{2N^{2}}{M},$ (16)
where we used the fact that $\binom{M}{N}/\binom{M+N-1}{N}>1-\frac{N^{2}}{M}$ and assumed $M\geq 2N^{2}$ in the last inequality for simplicity. Therefore, by Lemma 3, if we uniformly choose collision-free input (i.e., $\bm{t}$ in Eq. (1)), then the probability of observing collision outcomes is bounded by $\frac{2N^{2}}{M}$, for any unitary circuit.
Now, instead of using uniformly chosen input and fixed unitary circuit, we can rather fix the collision-free input $\bm{t}$ and draw random unitary circuits from $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ (see Definition 4 and Definition 5), which is the random circuit ensemble for the average-case hardness in the shallow-depth regime. Then, the random permutation circuit part from $\mathcal{P}$ permutes the collision-free input $\bm{t}$ uniformly random, and effectively makes input uniform distribution over possible collision-free configurations $G_{M,N}$. Hence, by choosing circuits from $\mathcal{H}_{\mathcal{A}}\mathcal{P}$, then the probability to observe collision outcomes is bounded by $\frac{2N^{2}}{M}$, and thus we have the following theorem.
Theorem 8 (Bosonic birthday paradox for shallow random circuits).
For a randomly chosen unitary circuit $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$, the probability to obtain collision outcomes is upper bounded by
$\operatorname*{\mathbb{E}}_{U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}}\left[\Pr_{\mathcal{D}_{U}}\left[\bm{s}\in B_{M,N}\right]\right]<\frac{2N^{2}}{M},$ (17)
where $\mathcal{D}_{U}$ is the output distribution according to $p_{\bm{s}}(U)$ defined in Eq. (1).
We have now obtained the desired result; the above theorem implies that for most of the random circuits over $\mathcal{H}_{\mathcal{A}}\mathcal{P}$, the sampler outputs a collision-free outcome with high probability in the dilute regime $M=\Omega(N^{2})$.
We lastly remark that this bosonic birthday paradox is not guaranteed solely by the local random circuit ensemble in Definition 4 itself, because of the absence of translation invariance property (in contrast to the global Haar random unitary circuit). Hence, the additional random permutation circuit at the input is crucial to guarantee the bosonic birthday paradox and suppress the occurrence of collision outcomes. Nevertheless, if one can prove that the local random circuit ensemble has a permutation-invariance property for a specific circuit architecture, then we can leave out this random permutation circuit part at the input. We leave in Appendix H numerical evidence of the bosonic birthday paradox for solely the local random circuit ensemble $\mathcal{H}_{\mathcal{A}}$ with our shallow-depth circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$.
Appendix C Proof of Theorem 3 (worst-case hardness)
To prove Theorem 3, we first introduce a crucial lemma, which states that any BQP circuit can be simulated by a constant-depth linear optical circuit under post-selection.
@@ Appendix C Proof of Theorem 3 (worst-case hardness)
For an arbitrary given poly-sized $n$-qubit quantum circuit $Q$, there exists a constant depth linear optical circuit $C_{0}$ such that for $M=\Omega(N)$ and $N={\rm{poly}}(n)$,
$|\bra{J}\hat{\mathcal{U}}(C_{0})\ket{I}|^{2}=c_{Q}|\bra{0}^{\otimes n}Q\ket{0}^{\otimes n}|^{2},$ (18)
where $c_{Q}=2^{-O(N)}$ is a $Q$-dependent constant which can be efficiently computed, $\hat{\mathcal{U}}(C_{0})$ is a unitary operator corresponding to the circuit $C_{0}$, and each of |$\ket{I}$ and |$\ket{J}$ is an $M$-mode Fock-state composed of $N$ single photon states and vacuum states for the rest modes.
@@
In this proof, we revise the main result in [33], incorporating a more quantitative analysis on the coefficient $c_{Q}$, which later determines the imprecision level allowed for the worst-case hardness of shallow-depth Boson Sampling. To do so, we first go through the main result in [33].
Given an arbitrary poly-sized quantum circuit $Q$ on $n$ qubits, there exists a measurement-based quantum computation (MBQC) scheme using constant depth brickwork graph state composed of maximally $m_{Q}=\text{poly}(n)$ number of qubits and $\Theta(m_{Q})$ number of CZ gates, with an explicit measurement basis sequence for the universal gate set {CX, T, H} to simulate $Q$ [51, 52, 53] (see, e.g., Fig. 4 (a)). Since the measurement outcomes for gate implementation are random, one needs adaptive measurements based on these measurement outcomes to implement $Q$ (which is indeed the essence of MBQC). However, if post-selecting desired measurement outcomes are allowed, one can rather post-select “good” outcomes that do not require any adaptive measurements to implement $Q$.
This brickwork qubit circuit can also be implemented in linear optical systems via the celebrated KLM scheme [54]. In their scheme, they use a dual-rail encoded state (|$\ket{01}$ or |$\ket{10}$ in the Fock basis) as a qubit, a beam splitter as a one qubit gate, and a 2-depth network of beam splitters for a probabilistic CZ gate which also requires additional two single photon ancillas for post-selecting successful CZ operation. Using this KLM scheme, Ref. [33] proposed a 4-depth linear optical circuit $C_{0}$, which implements the brickwork qubit state by post-selecting successful CZ operation (see Fig. 4 (b) or Fig. 3 in Ref. [33] for more details). Here, the required photon number is $N=\Theta(m_{Q})$ to implement the brickwork qubit state composed of $m_{Q}$ number of qubits and $\Theta(m_{Q})$ number of CZ gates, because each of qubit and CZ gate requires a constant number of photons in the KLM scheme. For more clarity, we illustrate in Fig. 4 the brickwork qubit circuit and corresponding linear optical circuit $C_{0}$ proposed by [33]. Also, the mode number is given by $M=\Omega(N)$, which is required for the dual-rail encoding in the KLM scheme. We remark that the mode number $M$ of $C_{0}$ can be arbitrarily enlarged by considering a circuit that “contains” the circuit in [33], which uses some of the modes for dual-rail encoding, leaving the rest of the modes as vacuum and adding only trivial (i.e., identity) gates at these unwanted modes. As the trivial gates leave the unwanted modes separated, this kind of mode number augmentation does not affect the post-selection probabilities.
To sum up, given an arbitrary quantum circuit $Q$, there exists a 4-depth linear optical circuit $C_{0}$ that can simulate the quantum circuit $Q$, by post-selecting successful CZ operation for all probabilistic CZ gates and post-selecting good measurement outcomes (that do not require any adaptive measurements) to implement each gate in the universal gate set {CX, T, H} composing the circuit $Q$. Here, as the explicit measurement basis sequence of brickwork qubit circuit for the universal gate set {CX, T, H} is given, one can find (i) proper coefficients of beam splitters in $C_{0}$ to implement the universal gate set via the KLM scheme and (ii) good measurement outcomes in the Fock basis. Hence, for the given quantum circuit $Q$, we can efficiently construct the constant-depth circuit $C_{0}$ (by means of determining its coefficients) and $M$-mode $N$-single photon Fock-states |$\ket{I}$ and |$\ket{J}$ such that Eq. (18) holds. Now, $c_{Q}$ becomes a post-selection probability to post-select (i) successful CZ operations for all CZ gates and (ii) good measurement outcomes to implement each gate in the universal gate set composing $Q$ via MBQC. In other words, $c_{Q}$ can be expressed as
@@
where $p_{k}$ denotes post-selection probability to implement $k$ gate (e.g., $p_{\text{CZ}}$ is $2/27$ in [55]), and $\Gamma_{k}$ denotes the number of $k$ gate to implement the circuit $Q$. Hence, given the post-selection probability $p_{k}$ for each gate, by counting the number of each gate to implement the circuit $Q$, $c_{Q}$ can be computed efficiently. To obtain the size of $c_{Q}$, note that we set the photon number $N=\Theta(m_{Q})$ to implement $m_{Q}$ qubit brickwork qubit circuit. As this $m_{Q}$ qubit brickwork qubit circuit can implement at most $O(m_{Q})$ number of gates, $Q$ has at most $O(m_{Q})=O(N)$ number of gates in the universal gate set {CX, T, H}. Also, the number of CZ gates is given as $O(N)$, and thus the total number of gates to implement $Q$ is at most $\sum_{k}\Gamma_{k}=O(N)$. Therefore, $c_{Q}$ has its amplitude $c_{Q}=2^{-O(N)}$, since all the post-selection probabilities $p_{k}$ are given as constants. This concludes the proof.
Next, we use the proposition by [42] which states that estimating the output probability of any BQP circuit is #P-hard for a certain additive imprecision.